Suppose that ABC Ltd is considering purchasing one of three new processing machines. Either machine would make it possible for the company to produce its products more efficiently. Estimates regarding each machine are provided below: Machine A Machine B Machine C Original cost $79,000 $110,000 $244,000 Estimated life 7 years 8 years 10 years Salvage value Nil Nil $30,000 Estimated annual cash inflows $30,000 $ 60,000 $58,500 Estimated annual cash outflows $ 7,000 $ 35,000 $18,500 A. If the projects cannot be repeated, which machine should ABC Ltd choose based on the NPV criteria at an 8% cost of capital? (9 marks) B. If the projects can be repeated, which machine should ABC Ltd choose based on the NPV criteria at an 8% cost of capital? (6 marks) C. Calculate the internal rate of return for Machine A? [Hint: internal rate of return is the rate which results in a zero NPV using linear interpolation], and discuss 1 drawback of the IRR against the NPV (5 marks)
1
Expert's answer
2020-06-03T18:59:49-0400

a) "NPVa=-79,000+(30,000-7,000)*\frac{(\frac{1}{1.08})^7-1}{\frac{1}{1.08}-1}=50,326.23"

"NPVb=-110,000+(60,000-35,000)*\frac{(\frac{1}{1.08})^8-1}{\frac{1}{1.08}-1}=45,159.25"

"NPVc=-244,000+(58,500-18,500)*\frac{(\frac{1}{1.08})^{10}-1}{\frac{1}{1.08}-1}-\frac{30,000}{1.08^{10}}=31,979.72"

 ABC Ltd should choose project A

b) For the comparable assessment, the same amount of time is needed. It is 14 years

A project will repeate 2 times; B project will repeate 1.75 times; C project will repeate 1.4 times

"NPVa=-79,000-\frac{79,000}{1.08^7}+(30,000-7,000)*\frac{(\frac{1}{1.08})^{14}-1}{\frac{1}{1.08}-1}=79,691.11"

"NPVb=-110,000-\frac{110,000*0.75}{1.08^8}+(60,000-35,000)*\frac{(\frac{1}{1.08})^{14}-1}{\frac{1}{1.08}-1}=68,022.22"

"NPVc=-244,000-\frac{244,000*0.4}{1.08^{10}}(58,500-18,500)*\frac{(\frac{1}{1.08})^{14}-1}{\frac{1}{1.08}-1}-\frac{30,000}{1.08^{14}}=56,738.48"

The project A still the most effective

c) "IRRa=-79,000+(30,000-7,000)*\frac{(\frac{1}{1+i})^7-1}{\frac{1}{1+i}-1}"

"0=-79,000+(30,000-7,000)*\frac{(\frac{1}{1+i})^7-1}{\frac{1}{1+i}-1}"

"i=0.3397=33.97\%"

IRR more difficult calculate